1.You have $39,000 in an account earning an interest rate of 4%. What are the equal beginning-of-month withdrawals you can make from this account before it is depleted in 22 years? Round to the nearest cent.

2.Your first job out of college will pay you $81,000 in year 1 (exactly one year from today). You estimate that your salary will grow at 5% per year. You plan to retire in 42 years (you'll receive 42 years of salary). If the applicable discount rate is 8%, what is the present value of these future earnings today? Round to the nearest cent.

please round answers to
the nearest cent thank you
Answer #1

1. PV =39000
Rate =4%/12
Number of Months =22*12 =264
The equal beginning of the month withdrawal =PV/(((1-(1+r)^-n)/r)*(1+r))=39000/(((1-(1+4%/12)^-264)/(4%/12*(1+4%/12))
=221.63

2. PV of future earnings =81000/1.08+81000*(1+5%)/1.08^2...…………+81000*(1+5%)^41/1.08^42
=81000/1.08*(1+1.05/1.08+1.05^2/1.08^2.....+1.05^41/1.08^41) =81000/1.08*(1-(1.05/1.08)^42)/(1-(1.05/1.08) =1872978.98